OpenStudy (falco276):

HELP!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!! The equation of a circle is shown below. (x -5)2 + (y + 2)2 = 64 The radius of the circle is ________ units. (No Multiple Choice)

OpenStudy (falco276):

@.Sam. @andriod09 @Gin_Ichimaru @ganeshie8 @Hunter15 @Creation @aaronq @lalokio

OpenStudy (falco276):

@hba @Frostbite @Hero @lexi14 @phi @Meriam

OpenStudy (falco276):

But what's the answer!?

OpenStudy (falco276):

I'm confused!!!!!!!!!!!!

OpenStudy (anonymous):

\[(x-h)^2+(y-k)^2=r^2\] is the equation of a circle with radius \(r\) and center \((h,k)\)

OpenStudy (anonymous):

in your case \(r^2=64\) what is \(r\) ?

OpenStudy (falco276):

And how do I translate that into the real equation?

OpenStudy (falco276):

I actuall don't know what R is. :/

OpenStudy (anonymous):

\[(x-h)^2+(y-k)^2=r^2\\(x-5)^2+(y+2)^2=64\] look almost identical

OpenStudy (falco276):

okay, so it's 64. Right?

OpenStudy (anonymous):

\(r^2\) is 64 but you are asked for the radius, which is \(r\) what number would you square to get 64?

OpenStudy (falco276):

8?

OpenStudy (anonymous):

yes

OpenStudy (falco276):

so that's the answer, right? 8?

OpenStudy (falco276):

hello? :(

OpenStudy (anonymous):

yes

OpenStudy (falco276):

okay, thanks. :)

OpenStudy (anonymous):

\(64=8^2\) so \(r=8\) yw

OpenStudy (falco276):

thanks. :)

OpenStudy (falco276):

Wait guys, I have another question. segment PR has an endpoint at (25, -5) and a midpoint of (18, -1). What is the value of the x-coordinate of the other endpoint?

OpenStudy (falco276):

@satellite73 @Loser66 @geoffb

OpenStudy (falco276):

(this one is also not multiple choice.)

OpenStudy (loser66):

post at a new post, friend